OpenStudy (anonymous):

How many grams of NH3 can be produced from 4.45mol of N2 and excess H2.

OpenStudy (mayankdevnani):

first write REACTION , \[\large \bf N_2+3H_2\rightarrow 2NH_3\] then,find LIMITING REAGENT. given,H2 is in excess ,so N2 is limiting. so find moles of NH3 with the help of LIMITING REAGENT i.e N2 \[\large \bf 1~mol~N_2 \rightarrow 2~mole~NH_3\] so, \[\large \bf 4.45~mol~N_2\rightarrow 4.45 \times 2~mol~NH_3=8.9~mol~NH_3\] now,we have moles of NH3 and molecular weight of NH3,now find weight of substance. \[\large \bf mole=\frac{weight}{molecular~weight}\] \[\large \bf weight=mole \times molecular~weight=8.9 \times 17=151.3\] so, \[\large \bf 151.3g~of ~NH_3~is~produced=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

hope you understand. @ipelayo1995

OpenStudy (gebooors):

Theoretically correct anwer. Synthesis of ammonia is an equilibrium reaction, so yield is not necessarily 100 %.